# Alphazero-like Tree-Search can guide large language model decoding and training

### v1

@@ Abstract
Large language models (LLMs) typically employ sampling or beam search, accompanied by prompts such as Chain-of-Thought (CoT), to boost reasoning and decoding ability. Recent work like Tree-of-Thought (ToT) and Reasoning via Planning (RAP) aim to augment the reasoning capabilities of LLMs by utilizing tree-search algorithms to guide multi-step reasoning. These methods mainly focus on LLMs’ reasoning ability during inference and heavily rely on human-designed prompts to activate LLM as a value function, which lacks general applicability and scalability. To address these limitations, we present an AlphaZero-like tree-search learning framework for LLMs (termed TS-LLM), systematically illustrating how tree-search with a learned value function can guide LLMs’ decoding ability. TS-LLM distinguishes itself in two key ways: (1) Leveraging a learned value function, our approach can be generally applied to different tasks beyond reasoning (such as RLHF alignment), and LLMs of any size, without prompting advanced, large-scale models. (2) It can guide LLM’s decoding during both inference and training. Empirical evaluations across reasoning, planning, and RLHF alignment tasks validate the effectiveness of TS-LLM, even on trees with a depth of 64.
@@ 1 Introduction
Pretrained autoregressively on extensive corpora, large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a) have demonstrated their potential in a wide range of natural language tasks. A plethora of recent studies have concentrated on improving LLMs task-solving capability, including curation of larger and higher-quality general or domain-specific data (Touvron et al., 2023a; Zhou et al., 2023; Gunasekar et al., 2023; Feng et al., 2023; Taylor et al., 2022), more sophisticated prompt design (Wei et al., 2022; Zhou et al., 2022; Creswell et al., 2022), or better training algorithms with Supervised Learning or Reinforcement Learning (RL) (Dong et al., 2023; Gulcehre et al., 2023; Rafailov et al., 2023). When training LLMs with RL, LLMs’ generation can be naturally formulated as a Markov Decision Process (MDP) and optimized with specific objectives. Following this formulation, ChatGPT (Ouyang et al., 2022) emerges as a notable success, optimizing LLMs to align human preference by leveraging RL from Human Feedback (RLHF) (Christiano et al., 2017).
Based on this formulation, LLMs’ generation can be further guided with planning algorithms such as tree search. The success of AlphaZero (Silver et al., 2017b) has strongly proved the effectiveness of tree search algorithms, especially Monte Carlo Tree Search (MCTS) (Kocsis & Szepesvári, 2006; Coulom, 2006), in solving large-scale MDPs, including Atari and Go. It attracts recent efforts on the potential of leveraging these principled planning approaches to improve LLMs’ reasoning ability.
Preliminary work in this field includes Tree-of-Thought (ToT) (Yao et al., 2023; Long, 2023) with depth/breadth-first search and Reasoning-via-Planing (RAP) (Hao et al., 2023) with MCTS. They successfully demonstrated a performance boost of searching on trees expanded by LLM through self-evaluation. Despite the strides, current methods come with distinct limitations. Firstly, these approaches mainly focus on enhancing LLMs’ reasoning ability, lacking general applicability in different kinds of tasks, such as RLHF alignment. Secondly, as a model-based algorithm, the crucial reward or value function of tree-search algorithms is mainly obtained by prompting LLMs for step-level or trajectory-level reasoning. As a result, such algorithms lack their applicability and heavily rely on both well-designed prompts and large-scale LMs, such as GPT4 or LLaMA-33B. Only LLMs in such scale can provide reliably innate critic ability for reward and value estimation.
@@
To address these problems, we introduce tree-search enhanced LLM (TS-LLM), an AlphaZero-like framework that utilizes tree-search to improve LLMs’ performance on general natural language tasks. TS-LLM extends previous work to AlphaZero-like deep tree-search with a learned value function and can guide LLM both during inference and training. Compared with previous work, TS-LLM has the following two new features:
@@
  • TS-LLM offers a generally applicable and scalable pipeline. It is generally applicable: (1) TS-LLM can be versatile to different tasks beyond reasoning. It can even solve RLHF alignment task in our experiment. (2) With a learned value function, TS-LLM can be applied to LLMs of any size. It does not require any well-designed prompts or advanced, large-scale LLMs. Our experiments show that TS-LLM can work for LLMs ranging from as small as 125M to as large as 7B. TS-LLM is also scalable: TS-LLM can conduct deep tree search, extending tree-search for LLM generation up to a depth of 64. This is far beyond 10 in ToT and 7 in RAP.
@@
  • TS-LLM can potentially serve as a new LLM training paradigm beyond inference decoding. By treating the tree-search operation as a policy improvement operator, we can conduct iterative processes of policy improvement and evaluation, to further train the LLM.
@@
Furthermore, we present an in-depth analysis of the core design elements in TS-LLM, delving into the features, advantages, and limitations of different variations. We also offer a novel and reasonable evaluation metric to fairly compare tree-search algorithms with other baselines. Through comprehensive empirical evaluations on reasoning, planning, and RLHF alignment tasks, we benchmark the performance of various TS-LLM  algorithm variants and underscore their superior capabilities. This showcases TS-LLM’s potential as a universal framework to guide LLM decoding and training.
@@ 2 Related work
Search guided reasoning in LLMs While most CoT approaches have used a linear reasoning structure, recent efforts have been made to investigate non-linear reasoning structures such as trees (Jung et al., 2022; Zhu et al., 2022). More recently, various approaches for searching on trees have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding and training framework with a learned value function, which is more generally applicable to both reasoning tasks and other scenarios like RLHF alignment tasks. Due to the limit of space, we leave a comprehensive comparison in Appendix A.
Finetuning LLMs with Augmentation Recent efforts have also been made to improve LLMs with augmented data. Rejection sampling is a simple and effective approach for finetuning data augmentation to improve LLMs’ ability on single/multiple task(s) such as multistep reasoning (Yuan et al., 2023a; Zelikman et al., 2022) and alignment with human preference (Dong et al., 2023; Bai et al., 2022). Given an augmented dataset, reinforcement learning approaches have been also used to finetune LLMs (Gulcehre et al., 2023; Luo et al., 2023). Compared to previous works, TS-LLM leverages tree search as a policy improvement operator to generate augmented samples to train both the LLMs and the value function.
Figure 1: Left: Sentence/Token-level node on Game24. Right: Overall framework: TS-LLM is a generally applicable and scalable pipeline with multiple choices of tree search approaches and search aggregation methods. Firstly, with LLM finetuned on the task training set, TS-LLM generates trajectories to learn value/ORM. By treating tree search as a policy improvement operator, it iteratively augments the training set and further improves the LLM and value/ORM.
3 Enhancing Large language model with Tree Search
In this section, we propose a versatile tree-search framework TS-LLM for guiding LLMs decoding and training. TS-LLM is summarized in the right part of Fig.1, and we conduct a systematic and comprehensive discussion about its key components. We start with the problem formulation with different tree/node designs, illustrating how LLMs’ generation can be framed into a tree-search pipeline. Secondly, we show how tree-search can help with LLM inference, including choices of specific tree-search algorithms, value-function training techniques, and search aggregation methods. Building upon tree-search enhanced inference, we present how tree-search can serve as a general policy improvement operator for guiding LLMs training. Lastly, we discuss the tree-search algorithm’s computation burdens for a fair comparison.
@@ 3 Enhancing Large language model with Tree Search
For a given natural language task, we can define a reward function $R(y_{t}|\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1})$ as the task performance feedback for intermediate generation $y_{t}$ at timestep $t$. Due to the lack of large-scale and high-quality intermediate reward labels for general tasks, it is usually a sparse reward setting where any intermediate reward from the first $T-1$ timestep is zero except the last $T$-th step. A typical case can be RLHF alignment task, where LLM can only receive the reward signal after it completes the full generation. Following the same logic, $\mathbf{y}$ can also be viewed as a sequence of sentences.
@@
Given the problem formulation above, we successfully transfer the problem of better generation to optimization for higher cumulative reward. In this paper, we focus on how we can optimize it with tree-search algorithms. A specific natural language task typically predefines the state space (with language) and reward function (with task objective/metrics). What remains is the definition of action space, or in the context of tree-search algorithm, the action node design problem.
@@
Tree search algorithms have validated their effectiveness for different action spaces in traditional RL research, including discrete action space (Silver et al., 2017a; Schrittwieser et al., 2020) and continuous action space (Hubert et al., 2021). For conducting tree-search on LLMs, we categorize the action space design into the following two aspects:
@@
  • Sentence-level action nodes: For the tasks that have a step/sentence-level structure(e.g. chain-of-thought reasoning), it is natural to treat each thought as a sentence-level action node. Fig. 1 shows an example of sentence-level action nodes. This is also the technique adopted by ToT (Yao et al., 2023) and RAP (Hao et al., 2023). For each non-terminal node, the search tree is expanded by sampling $m$ possible subsequent intermediate steps and dropping the duplicated generations.
@@
  • Token-level action nodes: Analogous to tree-search in discrete action space MDP, we can treat each token as a discrete action for LLM policy and the tree search can be conducted in token-level. We demonstrate an example of a Token-level action node in Fig. 1. For those tasks in which intermediate steps are not explicitly defined(e.g. RLHF alignment), splitting an output sequence into tokens might be a good choice.
@@
Both action space designs have their own advantages and limitations. By splitting the generation into sentences, sentence-level action nodes provide a relatively shallow tree, simplifying the tree-search process. However, the large sample space of sentence-level generation makes full enumeration of all possible sentences infeasible. We have to subsample $k$ nodes during the expansion, similar to the idea of Sampled MuZero(Hubert et al., 2021). Such subsampling results in the gap, determined by $k$, between the tree-search space and the LLM generation space. In addition, sentence-level action node often brings extra computation burdens, which will be further discussed in Section 3.4. For token-level action nodes, though it can get rid of the search space discrepancy and extra computational burdens, it greatly increases the depth of the tree, making tree-search more challenging.
@@ 3.2 Guiding LLM inference decoding with Tree Search
One of the benefits of tree-search algorithms is that they can optimize the cumulative reward by mere search, without any gradient calculation or update. Tree-search algorithms can enhance LLM generation without any further finetuning. In this section, given a fixed LLM policy, we present the full pipeline to illustrate how to guide LLM inference with tree search approaches.
3.2.1 Learning a Value Function
As a model-based RL algorithm, tree search typically requires three models, including dynamic transition model $g$, reward model $\hat{r}$(Schrittwieser et al., 2020), and value function $v$ Silver et al. (2017a). The dynamic transition in the LLM generation, given state $s$ and action $a$, is known as: $g\big{(}s=(\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1}),a=y_{t}\big{)}=(\mathbf{x}_{0:L-1},\mathbf{y}_{0:t})$. The value function $v$ and reward model $\hat{r}$ are the main issues. ToT and RAP obtain these two models by prompting advanced LLMs, such as GPT-4 or LLaMA-33B, because only LLMs of such scale can serve as sufficiently reliable models.
To make the tree search algorithm generally applicable, our method leverages a learned value function $v_{\theta}(s)$ conditioned on state $s$ and a learned final-step outcome reward model (ORM) $\hat{r}_{\theta}$ (Uesato et al., 2022) since most tasks can be formulated as sparse-reward problems. We utilize a shared value network and reward model, whose structure is a decoder-only transformer with an MLP to output a scalar on each position of the input tokens. For a sentence-level expanded intermediate step $s_{t}$, we use the prediction scalar at the last token as its value prediction $v_{\theta}(s_{t})$. The final reward can be obtained at the last token when feeding the full sentences $(\mathbf{x}_{0:L-1},\mathbf{y}_{0:T-1})$ into the model.
Therefore, we use language model $\pi_{\theta}$ as the policy to sample generations using the task training dataset. With true label or a given reward function in training data, a set of sampled tuple $\mathcal{D}_{train}=\{(\mathbf{x}^{k},\mathbf{y}^{k},r^{k})\}_{k=1}^{N}$ of size $N$ can be obtained, where $\mathbf{x}^{k}$ is the input text, $\mathbf{y}^{k}=s^{k}_{0:T^{k}-1}$ is the output text of $T^{k}$ steps and $r^{k}=R(\mathbf{y}^{k}|\mathbf{x}^{k})$ is the ground-truth reward. Similar to the critic training in most RL algorithms, we construct the value target $z^{k}_{t}$ by TD-$\lambda$ or MC estimate on each single step $t$. The value network is optimized by mean squared error:
$L(\theta)=\displaystyle\mathbb{E}_{\mathcal{D}}\left[\sum\limits_{t=0}^{T^{k}-1}||v_{\theta}(s^{k}_{0:t}|\mathbf{x}^{k})-z^{k}_{t}||_{2}^{2}\right].$ (1)
The ORM $\hat{r}_{\theta}(\mathbf{y}_{0:T-1}|\mathbf{x}_{0:L-1})$ is learned with the same objective. Training an accurate value function and ORM is quite crucial for the tree-search process as they provide the main guidance during the search. We will further illustrate how different training conditions influence the learned value function, ORM, and corresponding tree-search performance in our experiment section.

### v2

@@ Abstract
Recent works like Tree-of-Thought (ToT) and Reasoning via Planning (RAP) aim to augment the reasoning capabilities of LLMs by using tree-search algorithms to guide multi-step reasoning. These methods rely on prompting a pre-trained model to serve as a value function and focus on problems with low search depth. As a result, these methods will not work in domains where the pre-trained LLM does not have enough knowledge to serve as an effective value function or in domains that require long-horizon planning. To address these limitations, we present an AlphaZero-like tree-search learning framework for LLMs (termed TS-LLM), systematically illustrating how tree-search with a learned value function can guide LLM decoding. TS-LLM distinguishes itself in two key ways. (1) Leveraging a learned value function and AlphaZero-like algorithms, our approach can be generally adaptable to a wide range of tasks, language models of any size, and tasks of varying search depths. (2) Our approach can guide LLMs during both inference and training, iteratively improving the LLM. Empirical results across reasoning, planning, alignment, and decision-making tasks show that TS-LLM outperforms existing approaches and can handle trees with a depth of 64.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023a) have demonstrated their potential in a wide range of natural language tasks. A plethora of recent studies have concentrated on improving LLMs task-solving capability, including curation of larger and higher-quality general or domain-specific data (Touvron et al., 2023a; Zhou et al., 2023; Gunasekar et al., 2023; Feng et al., 2023; Taylor et al., 2022), more sophisticated prompt design (Wei et al., 2022; Zhou et al., 2022; Creswell et al., 2022), or better training algorithms with Supervised Learning or Reinforcement Learning (RL) (Dong et al., 2023; Gulcehre et al., 2023; Rafailov et al., 2023). When training LLMs with RL, LLMs’ generation can be naturally formulated as a Markov Decision Process (MDP) and optimized with specific objectives. Following this formulation, ChatGPT (Ouyang et al., 2022) emerges as a notable success, optimizing LLMs to align human preference by leveraging RL from Human Feedback (RLHF) (Christiano et al., 2017).
LLMs can be further guided with planning algorithms such as tree search. Preliminary work in this field includes Tree-of-Thought (ToT) (Yao et al., 2023; Long, 2023) with depth/breadth-first search and Reasoning-via-Planing (RAP) (Hao et al., 2023) with MCTS. They successfully demonstrated a performance boost of searching on trees expanded by LLM through self-evaluation.
Figure 1: (a) Left: Two node expansion paradigms on Game24: sentence-level and token-level. We adopt sentence-level setting in this task. (b) Right: TS-LLM consists of an iterative process over tree-search and training. First, TS-LLM enhances LLM inference by tree-search to obtain improved trajectories, augmenting the training set. LLM can be further trained to improve by conducting policy distillation and value function learning (policy evaluation) over the augmented training set.
Despite these advances, current methods come with distinct limitations. First, the value functions in the tree-search algorithms are obtained by prompting LLMs. As a result, such algorithms lack general applicability and heavily rely on both well-designed prompts and the robust capabilities of advanced LLMs. Beyond the model requirements, we will also show in Sec. 4.2.1 that such prompt-based self-evaluation is not always reliable. Second, ToT and RAP use BFS/DFS and MCTS for tree search, restricting their capabilities to relatively simple and shallow tasks. They are capped at a maximum depth of only 10 or 7, which is significantly less than the depth achieved by AlphaZero in chess or Go (Silver et al., 2017). As a result, ToT and RAP might struggle with complex problems that demand large analytical depths and longer-term planning horizons, decreasing their scalability.
@@
To address these problems, we introduce tree-search enhanced LLM (TS-LLM), an AlphaZero-like framework that utilizes tree-search to improve LLMs’ performance on general natural language tasks. TS-LLM extends previous work to AlphaZero-like deep tree-search with a learned LLM-based value function which can guide the LLM during both inference and training. Compared with previous work, TS-LLM has the following two new features:
@@
  • TS-LLM offers a generally applicable and scalable pipeline. It is generally applicable: With a learned value function, TS-LLM can be applied to various tasks and LLMs of any size. Our learned value function can be more reliable than the prompt-based counterpart and does not require any well-designed prompts or advanced, large-scale LLMs. Our experiments show that TS-LLM can work for LLMs ranging from 125M to 7B parameters, providing better evaluation even compared with GPT-3.5. TS-LLM is also scalable: TS-LLM can conduct deep tree search, extending tree-search for LLM generation up to a depth of 64. This is far beyond 10 in ToT and 7 in RAP.
@@
  • TS-LLM can potentially serve as a new LLM training paradigm beyond inference decoding. By treating the tree-search operation as a policy improvement operator, we can conduct an iterative processes of improving the policy via tree search and then improving the policy through distillation and the value function through the ground-truth training labels on the tree search trajectories.
@@
Through comprehensive empirical evaluations on reasoning, planning, alignment, and decision-making tasks, we present an in-depth analysis of the core design elements in TS-LLM, delving into the features, advantages, and limitations of different variations. This showcases TS-LLM’s potential as a universal framework to guide LLM decoding and training.
@@ 2 Related Work
Search-Guided Reasoning in LLMs While most CoT approaches have used a linear reasoning structure, recent efforts have been made to investigate non-linear reasoning structures such as trees (Jung et al., 2022; Zhu et al., 2023). More recently, various approaches for searching on trees have been applied to find better reasoning paths, e.g. beam search in Xie et al. (2023), depth-/breadth-first search in Yao et al. (2023) and Monte-Carlo Tree Search in (Hao et al., 2023). Compared with these methods, TS-LLM is a tree search guided LLM decoding and training framework with a learned value function, which is more generally applicable to both reasoning tasks and other scenarios like RLHF alignment tasks. Due to the limit of space, we leave a comprehensive comparison in Appendix A.
Finetuning LLMs with Augmentation Recent efforts have also been made to improve LLMs with augmented data. Rejection sampling is a simple and effective approach for finetuning data augmentation to improve LLMs’ ability on single/multiple task(s) such as multistep reasoning (Yuan et al., 2023a; Zelikman et al., 2022) and alignment with human preference (Dong et al., 2023; Bai et al., 2022). Given an augmented dataset, reinforcement learning approaches have also been used to finetune LLMs (Gulcehre et al., 2023; Luo et al., 2023). Compared to previous works, TS-LLM leverages tree search as a policy improvement operator to generate augmented samples to train both the LLMs and the value function.
3 Enhancing LLMs with Tree Search
In this section, we propose a versatile tree-search framework TS-LLM for guiding LLMs decoding and training. We conduct a systematic and comprehensive analysis of its key components. TS-LLM is summarized in Fig. 1.
@@ 3.1 Problem Formulation
For a given natural language task, we can define a reward function $R(y_{t}|\mathbf{x}_{0:L-1},\mathbf{y}_{0:t-1})$ as the task performance feedback for intermediate generation $y_{t}$ at timestep $t$. Due to the lack of large-scale and high-quality intermediate reward labels for general tasks, it is usually a sparse reward setting where any intermediate reward from the first $T-1$ timestep is zero except the last $T$-th step. A typical case can be RLHF alignment task, where LLM can receive the reward signal after it completes the full generation. Following the same logic, $\mathbf{y}$ can also be viewed as a sequence of sentences.
@@
Given the problem formulation above, we successfully transfer the problem of better generation to optimization for higher cumulative reward. In this paper, we focus on how we can optimize it with tree-search algorithms. A specific natural language task typically predefines the state space (with language) and reward function (with task objective/metrics). What remains is the definition of action space, or in the context of tree-search algorithm, the action node.
@@
Tree search algorithms have validated their effectiveness for different action spaces in traditional RL research, including discrete action space (Silver et al., 2017; Schrittwieser et al., 2020a) and continuous action space (Hubert et al., 2021). For tree-search on LLMs, we consider the following two action space designs, as shown in the left side of Fig. 1.
@@
Sentence-level action nodes: For the tasks that have a step/sentence-level structure(e.g. chain-of-thought reasoning), it is natural to treat each thought as a sentence-level action node. This is also the technique adopted by ToT (Yao et al., 2023) and RAP (Hao et al., 2023). For each non-terminal node, the search tree is expanded by sampling several possible subsequent intermediate steps and dropping the duplicated generations.
@@
Token-level action nodes: Analogous to tree-search in discrete action space MDP, we can treat each token as a discrete action for LLM policy and the tree search can be conducted in token-level. For those tasks in which intermediate steps are not explicitly defined(e.g. RLHF), splitting an output sequence into tokens might be a good choice.
@@
Typically, the search space is determined by two algorithm-agnostic parameters, the tree max width $w$ and tree max depth $d$. In LLM generation, both action space designs have their advantages and limitations over the search space. By splitting the generation into sentences, sentence-level action nodes provide a relatively shallow tree (low tree max-depth), simplifying the tree-search process. However, the large sample space of sentence-level generation makes full enumeration of all possible sentences infeasible. We have to set a maximum tree width $w$ to subsample $w$ nodes during the expansion, similar to the idea of Sampled MuZero (Hubert et al., 2021) (The node will be fixed once it is expanded). Such subsampling results in the gap, determined by $w$, between the tree-search space and the LLM generation space. For token-level action nodes, though it can get rid of the search space discrepancy and extra computational burdens, it greatly increases the depth of the tree, making tree-search more challenging.
@@ 3.2 Guiding LLM Inference Decoding with Tree Search
One of the benefits of tree-search algorithms is that they can optimize the cumulative reward by mere search, without any gradient calculation or update. In this section, given a fixed LLM policy, we present the full pipeline to illustrate how to guide LLM inference decoding with tree search approaches.
3.2.1 Learning an LLM-based Value Function
For tree-search algorithms, how to construct reliable value function $v$ and reward model $\hat{r}$ is the main issue. ToT and RAP obtain these two models by prompting advanced LLMs, such as GPT-4 or LLaMA-33B. To make the tree search algorithm generally applicable, our method leverages a learned LLM-based value function $v_{\phi}(s)$ conditioned on state $s$ and a learned final-step outcome reward model (ORM) $\hat{r}_{\phi}$ since most tasks can be formulated as sparse-reward problems (Uesato et al., 2022). Since we mainly deal with language-based tasks, we utilize a shared value network and reward model whose structure is a decoder-only transformer with an MLP to output a scalar on each position of the input tokens. And typically, LLM value’s decoder is adapted from original LLM policy $\pi_{\theta}$’s decoder, or the LLM value $v_{\phi}$ and policy $\pi_{\theta}$ can have a shared decoder(Silver et al., 2017). For a sentence-level expanded intermediate step $s_{t}$, we use the prediction scalar at the last token as its value prediction $v_{\phi}(s_{t})$. The final reward can be obtained at the last token when feeding the full sentences $(\mathbf{x}_{0:L-1},\mathbf{y}_{0:T-1})$ into the model.
Therefore, we use language model $\pi_{\theta}$ as the policy to sample generations using the task training dataset. With true label or a given reward function in training data, a set of sampled tuple $\mathcal{D}_{\text{train}}=\{(\mathbf{x}^{j},\mathbf{y}^{j},r^{j})\}_{j}$ of size $|\mathcal{D}_{\text{train}}|$ can be obtained, where $\mathbf{x}^{j}$ is the input text, $\mathbf{y}^{j}=s^{j}_{0:T^{j}-1}$ is the output text of $T^{j}$ steps and $r^{j}=R(\mathbf{y}^{j}|\mathbf{x}^{j})$ is the ground-truth reward. Similar to the critic training in most RL algorithms, we construct the value target $z^{j}_{t}$ by TD-$\lambda$ (Sutton, 1988) or MC estimate (Sutton & Barto, 2018) on each single step $t$ The value network is optimized by mean squared error:
$L(\phi)=\displaystyle\text{E}_{\mathcal{D}}\left[\sum\limits_{t=0}^{T^{j}-1}||v_{\phi}(s^{j}_{0:t}|\mathbf{x}^{j})-z^{j}_{t}||_{2}^{2}\right].$ (1)
The ORM $\hat{r}_{\phi}(\mathbf{y}_{0:T-1}|\mathbf{x}_{0:L-1})$ is learned with the same objective. Training an accurate value function and ORM is quite crucial for the tree-search process as they provide the main guidance. We will further illustrate how to learn a reliable value function and ORM in our experiment section.